Circular No. 7315

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

HD 209458 In response to a request from the Central Bureau, D. W. Latham,

Harvard-Smithsonian Center for Astrophysics, communicates on behalf also of D. Charbonneau, T. M. Brown, M. Mayor, and T. Mazeh: “We have prepared an updated ephemeris for the transits of HD 209458. This ephemeris is based on 150 velocity observations dating back to Aug. 1997 (made with HIRES on Keck I, as part of the G Dwarf Planet Search under NASA time, with ELODIE on the 1.93-m telescope at Observatoire de Haute Provence, and with CORALIE on the new Swiss 1.2-m telescope at La Silla; a paper describing the spectroscopic results is in preparation) and on two complete transit observations obtained on 1999 Sept. 8 and 15 (made by Charbonneau and Brown with the STARE instrument at the High Altitude Observatory, National Center for Atmospheric Research, Boulder, and reported in a paper submitted to Ap.J. Lett. by Charbonneau, Brown, Latham, and Mayor). G. Torres has carried out a simultaneous solution of the radial velocities and the transit observations and provided the following orbital period and epoch of transit center: P = 3.52433 ± 0.00027 days, T_{c} = 2451430.8238 ± 0.0033 (HJD). Predictions for the center of the transits through the end of Dec. 1999 are: HJD 2451508.3590 ± 0.0080, 2451511.8833 ± 0.0083, 2451515.4077 ± 0.0086, 2451518.9320 ± 0.0089, 2451522.4563 ± 0.0091, 2451525.9806 ± 0.0094, 2451529.5050 ± 0.0097, 2451533.0293 ± 0.010, 2451536.554 ± 0.010, 2451540.078 ± 0.011, 2451543.602 ± 0.011. Observers should plan to monitor the star for at least 2 hr before and after these times of transit center. Our identification of HD 209458 as a prime target for transit observations was made possible by the many contributions of the G Dwarf Planet Search, ELODIE, and CORALIE teams: J. L. Beuzit, M. Burnet, G. A. Druckier, D. Naef, F. Pepe, C. Perrier, D. Queloz, N. Santos, J. P. Sivan, G. Torres, S. Udry, and S. Zucker.”

COMET C/1999 H1 (LEE) Total visual magnitude estimates: Oct. 18.96 UT, 9.4 (P. Morel, Cam-

brai, France, 0.20-m reflector); 30.83, 10.4 (K. Hornoch, Lelekovice, Czech Republic, 0.35-m reflector); Nov. 8.90, 10.8 (J. Carvajal, Avila, Spain, 0.32m reflector); 14.95, 11.8 (R. J. Bouma, Poio, Portugal, 0.30-m reflector).

1999 November 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green